OpenStudy (anonymous):

WILL MEDAL AND FAN What are the zeros of the function? f(x) = x^2 – 16x + 48 Choose all answers that are correct. Question 3 options: –12 –4 4 12

OpenStudy (ineedhelplz):

(x - 8)² = 16 (x - 8) = ± 4 x = 12 , x = 4 OR (x - 4)(x - 12) = 0 x = 4 , x = 12

OpenStudy (anonymous):

thank you!!

OpenStudy (ineedhelplz):

You're welcome :)